# On Vanishing Gradients, Over-Smoothing, and Over-Squashing in GNNs: Bridging Recurrent and Graph Learning

### v1

@@ Appendix D Additional empirical results
D.1 Additional MPNN Jacobians
Here, we present in Figure 6 the eigendistribution of the layerwise Jacobians of GCN, GIN (Xu et al., 2018) and Gated-GCN (Bresson & Laurent, 2017). Across the board, we observe similar contraction effects in the Jacobian as those presented in the main paper, with a long number of eigenvalues accumulating at zero, with no significant changes in the distribution during training. However, the maximum eigenvalues for both GIN and Gated-GCN are much larger than those of GCN. We also compare a nonlinear feedforward network and a nonlinear GCN in Figure 7.
Figure 6: Eigenvalues of layer-to-layer Jacobian of different GNN models.
Figure 7: Histogram of eigenvalue modulus of the layerwise Jacobian for a nonlinear convolutional and a nonlinear feedforward layer.
D.2 Over-smoothing
Here, we include additional results related to over-smoothing experiments. Figure 8 shows the effect of $||\Lambda||_{2}$ in GCN-SSM on different graph structures, showing that lower Jacobian norms leads to a rapid decay of the Dirichlet energy, whereas values closer to one result in a more stable energy evolution. This result is also confirmed by Figure 10 and Figure 11. The former presents the vectorized Jacobian for ADGN (Gravina et al., 2023), SWAN (Gravina et al., 2025), and PHDGN (Heilig et al., 2025) on Cora, while the latter the Dirichlet energy evolution of different models on different topologies. Notably, in Figure 11, ADGN, SWAN, and PHDGN exhibit stable Dirichlet energy across layers, and Figure 10 reveals that these Jacobian norms are close to one. These results confirm that stable dynamics also ensure a non-decaying Dirichlet energy, effectively preventing over-smoothing.
Figure 8: Dirichlet Energy evolution of GCN-SSM for different $||\Lambda||_{2}$ on different graph topologies. Left: Cora. Middle: Grid graph. Right: Texas.
Figure 9: Dirichlet Energy evolution of GAT-SSM for different $||\Lambda||_{2}$ on different graph topologies. Left: Cora. Middle: Grid graph. Right: Texas.
Figure 10: Vectorized Jacobian for ADGN (Gravina et al., 2023), SWAN (Gravina et al., 2025), and PHDGN (Heilig et al., 2025) on Cora. Left: ADGN. Middle: SWAN. Right: PHDGN.
@@ D.2 Over-smoothing
D.3 Link between delay and vanishing gradients
Here, we show how the delay term in (Gutteridge et al., 2023) is directly related to preventing vanishing gradients. We do so by showing that adding the delay term to a GCN is effective at preventing over-smoothing, see Figure 12, as well as by checking the histogram of eigenvalues of the Jacobian, see Figure 14.
@@ D.3 Link between delay and vanishing gradients
Figure 13: Left: Performance on the RingTransfer task for DRew (Gutteridge et al., 2023). Right: Effect of dissipativity on performance.
@@
D.4 Graph Property Prediction
Edge-of-chaos behavior and long-range propagation. To further support our claim that mitigating gradient vanishing is key to strong long-range performance, Figure 15 shows each method’s average Jacobian eigenvalue distance to the edge-of-chaos (EoC) region. The figure demonstrates that methods such as ADGN (Gravina et al., 2023) and SWAN (Gravina et al., 2025), which remain closer to EoC, effectively propagate information over large graph radii, resulting in superior performance across all three tasks. Figure 16 presents an ablation study on multiple ADGN variants, controlled by the hyperparameter $\gamma$, which governs the positioning of the Jacobian eigenvalues ($\gamma<0$ places them outside the stability region, $\gamma>0$ inside, and $\gamma=0$ on the unit circle). Notably, regardless of the initial value of $\gamma$, ADGN consistently converges towards the EoC region as performance improves.
@@ D.4 Graph Property Prediction
Complete results. Table 4 compares our method on graph property prediction tasks against a range of state-of-the-art approaches, including GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2018), GraphSAGE (Hamilton et al., 2017), GIN (Xu et al., 2018), GCNII (Chen et al., 2020), DGC (Wang et al., 2021), GRAND (Chamberlain et al., 2021b), GraphCON (Rusch et al., 2022), ADGN (Gravina et al., 2023), SWAN (Gravina et al., 2025), PH-DGN (Heilig et al., 2025), and DRew (Gutteridge et al., 2023). Our method achieves exceptional results across all three tasks, consistently surpassing MPNN baselines, differential equation-inspired GNNs, and multi-hop GNNs. These findings underscore how combining powerful model dynamics with improved connectivity provides substantial benefits in tasks that require long-range information propagation.
Table 4: Mean test set $log_{10}(\mathrm{MSE})$($\downarrow$) and std averaged on 4 random weight initializations on Graph Property Prediction tasks. The lower, the better. Baseline results are reported from (Gravina et al., 2023, 2025; Heilig et al., 2025).
@@
    + $\text{eig}(\Lambda)\approx 1$2.4442±0.0984 -3.5928±0.1026 -2.2583±0.0085
@@
D.5 Additional comments on LRGB tasks
@@
In our experiments with the LRGB tasks, we observe that the peptides-func task exhibits significantly longer-range dependencies than the peptides-struct task. Notably, the peptides-struct task performs best when the model is not initialized at the edge of chaos and requires fewer layers. Conversely, on peptides-struct the model performs best when it is set to be at the edge of chaos, and shows a monotonic performance increase with additional layers, with optimal results achieved when using forty layers.
@@
Furthermore, we highlight that while our experiments with a small hidden dimension adhere to the parameter budget established in (Dwivedi et al., 2022), increasing the hidden dimension ($d\uparrow$) to 256 causes us to exceed the 500k parameter budget limit, even though our model maintains the same number of parameters as a regular GCN. While this budget is a useful tool to benchmark different models, we highlight that this restriction results in models running with fewer layers and small hidden dimensions. However, a large number of layers is crucial for effective long-range learning in graphs that are not highly connected, while increasing the hidden dimension also directly affects the bound in Theorem 5.1. As such, we believe that this parameter budget indirectly benefits models with higher connectivity graphs, inadvertently hindering models that do not perform edge addition.
Appendix E Supplementary Related Work
Long-range propagation on GNNs.
@@ Appendix E Supplementary Related Work
Learning long-range dependencies on graphs involves effectively propagating and preserving information across distant nodes. Despite recent advancements, ensuring effective long-range communication between nodes remains an open problem (Shi et al., 2023). Several techniques have been proposed to address this issue, including graph rewiring methods, such as (Gasteiger et al., 2019; Topping et al., 2021; Karhadkar et al., 2022; Barbero et al., 2023; Gutteridge et al., 2023; Black et al., 2023), which modify the graph topology to enhance connectivity and facilitate information flow. Similarly, Graph Transformers enhance the connectivity to capture both local and global interactions, as demonstrated by (Ying et al., 2021; Dwivedi & Bresson, 2021; Shi et al., 2021; Kreuzer et al., 2021; Rampášek et al., 2022; Wu et al., 2023). Other approaches incorporate non-local dynamics by using a fractional power of the graph shift operator (Maskey et al., 2024), leverage quantum diffusion kernels (Markovich, 2023), regularize the model’s weight space (Gravina et al., 2023, 2025), or exploit port-hamiltonian dynamics (Heilig et al., 2025).
@@
One of the primary challenges in training recurrent neural networks lies in the vanishing (and sometimes exploding) gradient problem, which can hinder the model’s ability to learn and retain information over long sequences. In response, researchers have proposed numerous architectures aimed at preserving or enhancing gradients through time. Examples include Unitary RNNs (Arjovsky et al., 2016), Orthogonal RNNs (Henaff et al., 2016), coRNNs (Rusch & Mishra, 2020), Linear Recurrent Units (Orvieto et al., 2023), and Structured State Space Models (Gu et al., 2021; Gu & Dao, 2023). By leveraging properties such as orthogonality, carefully designed parameterizations, or alternative update mechanisms, these models seek to alleviate gradient decay and capture longer-range temporal relationships more effectively.
Dynamical-systems-inspired neural networks.
Since the introduction of Neural ODEs in Chen et al. (2018), there have been various methods that employ ideas of dynamical systems within neural networks, including continuous-time methods (Rubanova et al., 2019; Norcliffe et al., 2020; Calvo-Ordonez et al., 2023, 2024; Bergna et al., 2024; Moreno-Pino et al., 2024; Calvo-Ordoñez et al., 2025) or state-space approaches (Chang et al., 2023; Duran-Martin et al., 2024a, b). Within graph neural networks, we highlight PDE-GCN (Eliasof et al., 2021), GRAND (Chamberlain et al., 2021b), BLEND (Chamberlain et al., 2021a) and Neural Sheaf Diffusion (Bodnar et al., 2022).

### v2

@@ Appendix D Additional empirical results
D.1 Additional Over-Smoothing Results
Here, we include additional results related to over-smoothing experiments. Figure 6 shows the effect of $\Lambda||_{2}$ in GCN-SSM on different graph structures (similarly Figure˜7 for GAT-SSM), showing that lower Jacobian norms leads to a rapid decay of the Dirichlet energy, whereas values closer to one result in a more stable energy evolution. This result is also confirmed by Figure 8 and Figure 9. The former presents the vectorized Jacobian for ADGN [47], SWAN [48], and PHDGN [59] on Cora, while the latter the Dirichlet energy evolution of different models on different topologies. Notably, in Figure 9, ADGN, SWAN, and PHDGN exhibit stable Dirichlet energy across layers, and Figure 8 reveals that these Jacobian norms are close to one. These results confirm that stable dynamics also ensure a non-decaying Dirichlet energy, effectively preventing over-smoothing.
Figure 6: Dirichlet Energy evolution of GCN-SSM for different $\Lambda||_{2}$ on different graph topologies. Left: Cora. Middle: Grid graph. Right: Texas.
Figure 7: Dirichlet Energy evolution of GAT-SSM for different $\Lambda||_{2}$ on different graph topologies. Left: Cora. Middle: Grid graph. Right: Texas.
Figure 8: Vectorized Jacobian for ADGN [47], SWAN [48], and PHDGN [59] on Cora. Left: ADGN. Middle: SWAN. Right: PHDGN.
@@ D.1 Additional Over-Smoothing Results
D.2 Link between delay and vanishing gradients
Here, we show how the delay term in [55] is directly related to preventing vanishing gradients. We do so by showing that adding the delay term to a GCN is effective at preventing over-smoothing, see Figure 10, as well as by checking the histogram of eigenvalues of the Jacobian, see Figure 12. Figure˜11 shows the effect of $\Lambda||_{2}$ in DRew-SSM on the performance on the RingTransfer task, showing that lower Jacobian norms leads to a rapid performance decay, i.e., poor long-range propagation.
@@ D.2 Link between delay and vanishing gradients
Figure 11: Left: Performance on the RingTransfer task for DRew [55]. Right: Effect of dissipativity on performance.
@@
D.3 Graph Property Prediction
Edge-of-chaos behavior and long-range propagation. To further support our claim that mitigating gradient vanishing is key to strong long-range performance, Figure 13 shows each method’s average Jacobian eigenvalue distance to the edge-of-chaos (EoC) region. The figure demonstrates that methods such as ADGN [47] and SWAN [48], which remain closer to EoC, effectively propagate information over large graph radii, resulting in superior performance across all three tasks. Figure 14 presents an ablation study on multiple ADGN variants, controlled by the hyperparameter $\gamma$, which governs the positioning of the Jacobian eigenvalues ($\gamma<0$ places them outside the stability region, $\gamma>0$ inside, and $\gamma=0$ on the unit circle). Notably, regardless of the initial value of $\gamma$, ADGN consistently converges towards the EoC region as performance improves.
@@ D.3 Graph Property Prediction
Complete results. Table 4 compares our method on graph property prediction tasks against a range of state-of-the-art approaches, including GCN [67], GAT [101], GraphSAGE [56], GIN [106], GCNII [26], DGC [103], GRAND [23], GraphCON [91], ADGN [47], SWAN [48], PH-DGN [59], and DRew [55]. Our method achieves exceptional results across all three tasks, consistently surpassing MPNN baselines, differential equation-inspired GNNs, and multi-hop GNNs. These findings underscore how combining powerful model dynamics with improved connectivity provides substantial benefits in tasks that require long-range information propagation.
Table 4: Mean test set $log_{10}(\mathrm{MSE})$($\downarrow$) and std averaged on 4 random weight initializations on Graph Property Prediction tasks. The lower, the better. Baseline results are reported from [47, 48, 59].
@@
   + $\text{eig}(\Lambda)\approx 1$ -2.4442±0.0984 -3.5928±0.1026 -2.2583±0.0085
@@
D.4 Additional comments on LRGB tasks
@@
In our experiments with the LRGB tasks, we observe that the peptides-func task exhibits significantly longer-range dependencies than the peptides-struct task. Notably, the peptides-struct task performs best when the model is not initialized at the edge of chaos and requires fewer layers. Conversely, on peptides-func the model performs best when it is set to be at the edge of chaos, and shows a monotonic performance increase with additional layers, with optimal results achieved when using forty layers.
@@
Furthermore, we highlight that while our experiments with a small hidden dimension adhere to the parameter budget established in [36], increasing the hidden dimension ($d\uparrow$) to 256 causes us to exceed the 500k parameter budget limit, even though our model maintains the same number of parameters as a regular GCN. While this budget is a useful tool to benchmark different models, we highlight that this restriction results in models running with fewer layers and small hidden dimensions. However, a large number of layers is crucial for effective long-range learning in graphs that are not highly connected, while increasing the hidden dimension also directly affects the bound in Theorem 5.1. As such, we believe that this parameter budget indirectly benefits models with higher connectivity graphs, inadvertently hindering models that do not perform edge addition.
To further strengthen the comparison on real-world tasks, in Table˜5 we report results against GRED [32], a method inspired by state-space models (SSMs). We note that our model achieves superior performance on peptides-func, while GRED performs better on peptide-struct. Although GRED shares a few conceptual similarities with our approach, we emphasize several key differences: (i) GRED aggregates information from multiple neighborhoods at each layer, whereas our model aggregates from a single-hop neighborhood per layer. (ii) GRED’s SSM update operates inward, beginning with distant nodes and moving toward the root node, whereas in our model each additional layer aggregates from progressively more distant neighbors, moving outward. This propagation pattern is more consistent with the standard MPNN framework. (iii) Our model employs a fixed and non-learned state matrix, removing the need for additional constraints to guarantee stable learning dynamics. (iv) Unlike GRED, we do not perform weight sharing when applying k-hop aggregation. Finally, we stress that the primary goal of our model design is not to achieve state-of-the-art performance (as is the case with GRED) but rather to construct a minimal and controllable framework that allows us to isolate and study specific phenomena of interest.
Table 5: Comparison with GRED on the LRGB dataset. $d\uparrow$ adds latent dims.
Model Pept-func Pept-struct
[…]
kGCN-SSM 69.02±0.22 28.98±0.32
$d\uparrow$ 72.12±0.27 27.01±0.07
GRED 70.85±0.27 25.03±0.19
D.5 Scalability Results
In terms of runtime, GNN-SSM was deliberately designed to retain the simplicity and efficiency of its backbone (e.g., GCN). Our formulation introduces only two additional (fixed) matrices to store w.r.t. its backbone, and involves one element-wise addition and two extra matrix multiplications per layer. These additions have a negligible impact on memory and runtime. Consequently, our model retains the complexity of its backbone while substantially improving performance, see Tables˜6 and 7 below.
Table 6: Epoch Time (sec.) for GCN and GCN-SSM when performing node classification of the Cora dataset.
Layers GCN GCN-SSM
5 0.009 0.009
10 0.015 0.017
20 0.025 0.031
30 0.041 0.046
40 0.053 0.051
50 0.066 0.075
60 0.078 0.089
Table 7: Accuracy on ogbn-arxiv[62].
Model ogbn-arxiv
MPNNs
  GAT 72.01±0.20
  GCN 70.84±0.23
  GraphSAGE 71.49±0.27
Graph Transformers
  NodeFormer 59.90±0.42
  GraphGPS 70.92±0.04
  GOAT 72.41±0.40
  EXPHORMER + GCN 72.44±0.28
  SPEXPHORMER 70.82±0.24
Our
  GCN-SSM 72.49±0.16
D.6 State-Space Matrices Sensitivity
Empirically, we have that sharing $\Lambda$ across layers did not alter performance: using a single fixed $\Lambda$ yielded the same accuracy as training each $\Lambda_{i}$ with identical dynamics. Fixing $\Lambda$ ensures the system remains at the edge of chaos during training. Preserving this prior under a trainable $\Lambda$ would require additional constraints, in line with stabilization techniques used in RNN architectures, see Table 8 below.
Table 8: Performance comparison for different design choices when performing node classification of the Cora dataset.
Layers GCN-SSM (shared) GCN-SSM (no sharing) GCN-SSM (trained $\boldsymbol{\Lambda}$)
5 76.3 78.3 71.3
10 78.5 78.5 71.1
20 81.2 77.8 49.9
30 78.1 79.6 33.8
40 77.5 78.5 31.9
50 76.4 74.6 31.9
60 77.8 77.4 31.9
D.7 Preliminary Results on Heterophilic benchmarks
To further complement the empirical evaluation of our method, we conducted experiments on three heterophilic tasks from [85]: Amazon Ratings, Roman Empire, and Minesweeper. We compared GCN-SSM against the original GCN results reported in [85], as well as a GCN with the same depth as our model (denoted GCN (Optimal L)). The depth for each task was tuned based on validation performance, resulting in optimal layer counts of 4 for Amazon Ratings, 9 for Roman Empire, and 12 for Minesweeper. Results, averaged over three random seeds, are reported in Table˜9 and show that GCN-SSM outperforms the original GCN by approximately 8 accuracy points on average, and exceeds GCN (Optimal L) by 31 accuracy points.
Table 9: Performance comparison of our GCN-SSM with respect to the original GCN results reported in [85], as well as a GCN with the same depth as our model (denoted GCN (Optimal L)) on the heterophilic datasets from [85].
Model Amazon Ratings Roman Empire Minesweeper
[…]
GCN ([85]) 47.70±0.63 73.69±0.74 89.75±0.52
GCN (Optimal L) 47.57±0.67 33.83±0.41 60.84±0.89
GCN-SSM (Optimal L) 51.72±0.33 88.37±0.60 96.02±0.52
Appendix E Additional Details and Comments on Over-Smoothing
E.1 Choice of Feature Distance Measure
Throughout the paper we adopt the unnormalized Dirichlet energy
$\mathcal{E}(\mathbf{H})\;=\;\operatorname{tr}\,\bigl(\mathbf{H}^{\top}\mathbf{L}\,\mathbf{H}\bigr)\;=\;\sum_{(u,v)\in\mathcal{E}}\|\mathbf{h}_{u}-\mathbf{h}_{v}\|_{2}^{2},$
This choice aligns with several well cited papers in the over‐smoothing literature [91, 90]. Moreover, the connection we unravel between vanishing gradients and over‐smoothing also explains why techniques borrowed from recurrent architectures [91, 47] are expected to mitigate feature collapse in GNNs.
While our theoretical analysis focuses on $\mathcal{E}(\mathbf{H})$ for mathematical simplicity, we will now also evaluate an alternative smoothness measure to ensure our insights generalize beyond this choice. In particular, we use the smoothness measure employed in other over-smoothing works [105, 95]
$\mu(\mathbf{H})=\bigl\|\mathbf{H}-\mathbf{1}\,\boldsymbol{\gamma}_{\mathbf{H}}\bigr\|_{F},\quad\boldsymbol{\gamma}_{\mathbf{H}}=\frac{\mathbf{1}^{\top}\mathbf{H}}{N},$
We report in Figure˜15 empirical experiments on this measure, which empirically shows that the qualitative trends predicted by our unnormalized‐energy theory also manifest under this alternative metric. Although formal equivalence between these energies and our collapse proofs is not explored, this empirical alignment provides strong justification for the broader applicability of our analysis to the broader literature on oversmoothing.
Figure 15: GCN-SSM on Cora with $\mu(\mathbf{H})$ smoothing measure.
E.2 The effect of the Jacobian spectrum on node classification performance
In order to assess how the spectral properties of the layer‐wise Jacobian influence node‐classification performance, we carried out the following two experiments on the Cora citation network.
First, we systematically varied the spectrum of the diagonal $\Lambda$ matrix in our GCN-SSM backbone. For each number of layers $n\in\{5,10,20,30,40,50,60\}$, we set the maximal eigenvalue of $\Lambda$ to one of $\{1,\,0.66,\,0.33\}$, retrained the model, and recorded the best test accuracy. As shown in Table 10, moving the spectrum of $\Lambda$ away from unity leads to a pronounced degradation in accuracy, indicating that keeping the Jacobian eigenvalues near one is crucial for stable and effective information propagation across many layers.
Second, to isolate the contribution of the SSM backbone itself, we performed an analogous spectrum‐shaping experiment directly on the weight matrix $\mathbf{W}$ of a vanilla GCN (i.e. without any SSM components). We scaled $\mathbf{W}$ so that its spectral radius lay near the edge of stability (spectral norm $\approx 1$), but otherwise left the model architecture and training unchanged. Despite matching the Jacobian stability regime, these “spectrally tuned” vanilla GCNs failed to achieve the accuracy improvements seen with the full GCN-SSM backbone (first column of Table 10). This confirms that merely tuning $\mathbf{W}$’s spectrum is insufficient: the structured state-space dynamics provided by the SSM backbone are essential for the observed performance gains.
Table 10: Node‐classification accuracy on Cora when varying the spectrum of the backbone Jacobian ($\Lambda$) and comparing to vanilla GCN models whose weight matrix $\mathbf{W}$ is spectrally tuned near the edge of stability.
$n_{\mathrm{layers}}$ $\text{eig}(\Lambda)=1$ $\text{eig}(\Lambda)=0.66$ $\text{eig}(\Lambda)=0.33$ $\text{eig}(W)=1$ (GCN)
5 81.30 78.10 74.00 71.90
10 78.70 61.90 56.00 33.80
20 78.90 48.00 30.20 31.80
30 80.00 39.70 18.00 22.30
40 77.90 34.60 20.30 25.60
50 77.70 29.10 24.90 23.80
60 77.70 20.50 20.40 29.10
E.3 On Residual Connections and Gating
Several prior works have employed residual connections in GNNs to counteract over-smoothing, for example, see [69]. In fact, these residual‐GNN designs can be viewed as a special case of our approach, corresponding to the choice $\boldsymbol{\Lambda}=\mathbf{I}$. Under this constraint, the model outperforms a standard, memoryless GCN (see Fig. 3), but only by accumulating node features in an unstructured way.
In order to guarantee stable propagation dynamics, the spectral properties of the propagation matrix $\mathbf{B}$ play an important role: by appropriately “damping” incoming signals, one can stabilize the system’s behavior and prevent feature collapse or explosion.
To highlight the role of $\mathbf{B}$, we conducted an ablation study on the Cora dataset in which the propagation matrix was removed from our method. Table˜11 reports the results, showing that for deeper networks, the inclusion of $\mathbf{B}$ consistently improves performance. This behavior can be understood theoretically from the perspective of the Jacobian. While the eigenvalue structure induced by places the system near the edge of stability (with eigenvalues close to 1 in modulus), it is B that controls how much these dynamics "disperse" around the critical point (1,0) shown in Figure˜2. In this setting, too much dispersion can lead to instability.
Table 11: Node‐classification accuracy on Cora with and without the propagation matrix $\mathbf{B}$ across different number of layers.
$n_{\mathrm{layers}}$ Without $\mathbf{B}$ With $\mathbf{B}$
2 80.2 79.6
4 77.8 80.4
8 80.0 79.8
16 74.1 79.8
32 31.9 79.9
64 31.9 79.9
128 14.9 77.9
256 13.0 80.2
Appendix F Supplementary Related Work and Limitations
Long-range propagation and depth GNNs.
@@ Appendix F Supplementary Related Work and Limitations
Learning long-range dependencies on graphs involves effectively propagating and preserving information across distant nodes. Despite recent advancements, ensuring effective long-range communication between nodes remains an open problem [96]. Several techniques have been proposed to address this issue, including graph rewiring methods, such as [42, 100, 65, 8, 55, 13], which modify the graph topology to enhance connectivity and facilitate information flow. Similarly, Graph Transformers enhance the connectivity to capture both local and global interactions, as demonstrated by [109, 35, 97, 68, 87, 104]. Other approaches incorporate non-local dynamics by using a fractional power of the graph shift operator [74], leverage quantum diffusion kernels [73], regularize the model’s weight space [47, 48, 49], exploit port-hamiltonian dynamics [59], or use a graph adaptive method based on a learnable ARMA framework [37]. Some methods which have increased the depth of GNNs include [69, 71]. Alternative methods combine the spatial and spectral perspective of graph propagation to enable long-range communication between nodes, either by using spatially and spectrally parametrized graph filters [43] or spectral filters based on Chebyshev polynomials [57]. Recent work has also proposed new datasets to test long range dependencies [70], as well as ways to theoretically measure them [4].
@@
One of the primary challenges in training recurrent neural networks lies in the vanishing (and sometimes exploding) gradient problem, which can hinder the model’s ability to learn and retain information over long sequences. In response, researchers have proposed numerous architectures aimed at preserving or enhancing gradients through time. Examples include Unitary RNNs [2], Orthogonal RNNs [60], Linear Recurrent Units [81], and Structured State Space Models [53, 52]. By leveraging properties such as orthogonality, carefully designed parameterizations, or alternative update mechanisms, these models seek to alleviate gradient decay and capture longer-range temporal relationships more effectively. We highlight recent work that also extends these insights to transformer based architectures [6].
Dynamical systems and physics inspired neural networks.
Since the introduction of Neural ODEs in [27], there have been various methods that employ ideas of dynamical systems within neural networks, including continuous-time methods [89, 78, 19, 20, 12, 77, 21] or state-space approaches [25, 33, 34, 37]. Within graph neural networks, we highlight PDE-GCN [38], GRAND [23], BLEND [22] and Neural Sheaf Diffusion [14] in the static graph domain, while CTAN [50] and TG-ODE [51] in the temporal graph domain [46]. Other approaches which leverage other type of physics-inspired inductive biases such as topological latent space modelling include [24, 54, 15, 92].
Broader Impact, Limitations and Future Work.
We believe our work opens up a number of interesting directions that aim to bridge the gap between graph and sequence modeling. In particular, we hope that this work will encourage researchers to adapt vanishing gradient mitigation methods from the sequence modeling community to GNNs, and conversely explore how graph learning ideas can be brought to recurrent models. In our work, we mostly focused on GCN and GAT type updates, but we believe that our analysis can be extended to understand how different choices of updates and non-linearities affect training dynamics, which we leave for future work.